### v2

# Fire [Uncaptioned image]: Fact-checking with Iterative Retrieval and Verification

Abstract
Fact-checking long-form text is challenging, and it is therefore common practice to break it down into multiple atomic claims. The typical approach to fact-checking these atomic claims involves retrieving a fixed number of pieces of evidence, followed by a verification step. However, this method is usually not cost-effective, as it underutilizes the verification model’s internal knowledge of the claim and fails to replicate the iterative reasoning process in human search strategies. To address these limitations, we propose Fire, a novel agent-based framework that integrates evidence retrieval and claim verification in an iterative manner. Specifically, Fire employs a unified mechanism to decide whether to provide a final answer or generate a subsequent search query, based on its confidence in the current judgment. We compare Fire with other strong fact-checking frameworks and find that it achieves slightly better performance while reducing large language model (LLM) costs by an average of 7.6 times and search costs by 16.5 times. These results indicate that Fire holds promise for application in large-scale fact-checking operations. Our code is available at https://github.com/mbzuai-nlp/fire.git.
Fire [Uncaptioned image]: Fact-checking with Iterative Retrieval and Verification
Zhuohan Xie1  Rui Xing1, 2  Yuxia Wang1  Jiahui Geng1 Hasan Iqbal1  Dhruv Sahnan1  Iryna Gurevych1  Preslav Nakov1 1MBZUAI, 2The University of Melbourne {zhuohan.xie, preslav.nakov}@mbzuai.ac.ae
1 Introduction
“Every man has a right to his opinion, but no man has a right to be wrong in his facts.” - Bernard M. Baruch
Large language models (LLMs) have demonstrated exceptional performance across a wide range of tasks, including both language comprehension and generation (Zhao et al., 2023; Xie et al., 2023a). Consequently, LLMs are now widely applied in various domains (Xie et al., 2023b), and many users increasingly rely on the information they provide. However, this reliance is problematic, as LLMs are capable of producing outputs that are highly confident but factually incorrect, highlighting the critical need for robust fact-checking systems (Akhtar et al., 2023). However, fact-checking the entire output of LLMs in a single step is highly challenging. To address this, Min et al. (2023) proposed decomposing the content into multiple atomic claims, each of which can be individually verified. While this approach simplifies the fact-checking process, assessing the veracity of these atomic claims remains complex, especially when many require sourcing evidence from the web. Indeed, identifying the most relevant evidence online is a key challenge in fact-checking pipelines (Wang et al., 2024a).
To address this issue, conventional methods, such as FacTool and Factcheck-GPT (Chern et al., 2023; Wang et al., 2024a), frame the problem as a question-answering task, as illustrated on the left side of Figure 1. In these approaches, an LLM is prompted to generate N relevant questions, which are then used as search queries by a web search tool. The search results serve as evidence for LLM to determine the factuality of the claim. However, we argue that this process is inefficient in two key aspects. First, it underutilizes the internal knowledge already embedded in LLMs during pre-training. For claims involving common knowledge or widely known events, the LLM could confidently assess the claim without relying on external information. Second, generating multiple search queries concurrently does not align with the typical human reasoning process during search (Hu et al., 2023). Humans tend to begin with an initial query, gather information, and then refine their perspective on the claim, which often leads to the formulation of more effective follow-up queries.
Figure 1: Comparisons between Fire and previous frameworks. Previous frameworks typically treat web search and claim verification as distinct processes. In contrast, Fire integrates interactive retrieval and verification.
To address this gap, we introduce Fact-checking with Iterative Retrieval and VErification (Fire), an innovative agent-based framework that integrates both the internal knowledge of LLMs and external knowledge sources by unifying the verification process and search query generation into a single step. As illustrated on the right side of Figure 1, Fire employs a mechanism to decide whether to produce the final answer or generate a new search query, continuing the evidence-seeking process. This decision is based on the model’s confidence in its judgment. The closest related work to us is Safe (Wei et al., 2024), depicted in the center of Figure 1. Their method generates web search queries iteratively and subsequently verifies whether the entire retrieved evidence supports the claim. However, this approach lacks flexibility, as it treats evidence retrieval and claim verification as distinct processes, requiring a predetermined fixed number of searches regardless of the claim’s complexity. In contrast, our approach integrates evidence retrieval and claim verification into an iterative framework, encouraging the language model to verify based on its own knowledge and conduct searches only when necessary. Our experiments demonstrate that our method significantly reduces the computational costs of LLMs by an average factor of 7.6, as well as search-related costs by a factor of 16.5, all while maintaining fact-checking performance.
In summary, our contributions are as follows:
  • We present Fire, a simple yet effective interactive framework for fact-checking. Through extensive experiments conducted across multiple datasets, we demonstrate that our framework significantly reduces the LLM computational and search costs, making it a better option for large-scale production.
  • Our ablation studies demonstrate that the step-by-step reasoning process enhances the model’s confidence in fact-checking, particularly with GPT-4o-mini. For GPT-4o, we observed a similar trend; however, the effect was not as pronounced as that seen with GPT-4o-mini.
  • We conducted an error analysis and identified several quality issues in the current benchmark datasets, including the presence of ungrounded claims. Additionally, the strict reasoning capabilities of the LLM may incorrectly classify some debatable claims as non-factual.
2 Related Work
LLM Factuality
Despite the remarkable capabilities of LLMs (Brown et al., 2020; OpenAI, 2023; Zhao et al., 2023), the auto-regressive learning objective does not inherently offer strong guarantee or enforce the learning of factual accuracy in the training process, making these models produce content that deviates from real-world facts (Wang et al., 2024b). On average, there are 5%-10% false claims in responses of GPT-4 (OpenAI, 2023) and LLaMA-2 (Touvron et al., 2023) on world-knowledge questions (Iqbal et al., 2024). Retrieval-augmented generation (Guu et al., 2020) and post-generation fact-checking are essential for ensuring accurate knowledge dissemination. Retrieving highly relevant information plays a pivotal role in both guiding generation as a reference and determining verification results in fact-checking systems (Wang et al., 2024a).
The retriever and verifier are the most resource-consuming components in fact-checking systems, in terms of time and cost. Even with the inexpensive APIs (e.g., Serper at 0.001 USD per request and GPT-3.5-turbo for verification), verifying an atomic claim costs approximately 0.02 USD, making extensive verification impractical for general users (Iqbal et al., 2024). This high cost limits the ability to verify large volumes of LLM responses, potentially contributing to the spread of misinformation. Our framework aims to minimize the costs in these two steps, enabling affordable verification for general users. This allows them to easily verify suspicious or doubtful information, enhancing the dissemination of factual information.
Fact Checking with Agents
The recent advancements in LLMs have spurred significant research on LLM-powered agents, which are capable of reasoning about their environment and making decisions by either invoking external tools or performing internal actions (Wang et al., 2024c). These agent frameworks typically consist of several components, including reasoning, tool usage, memory, and multi-agent debate (Masterman et al., 2024), many of which can be seamlessly integrated into fact-checking pipelines to enhance the performance of traditional fact-checking systems. For example, recent works have endowed systems with the ability to call external tools, such as search engines (Chern et al., 2023; Wang et al., 2024a; Wei et al., 2024; Cheng et al., 2024), recognizing that many claims in the field require additional information for verification. During the verification stage, Sun et al. (2024) proposed a Markov Chain-based multi-agent debate approach to ensure more rigorous verification by enabling collaborative decision-making among agents based on retrieved evidence. Our work differs from previous approaches by combining the evidence retrieval and verification stages, leveraging agents’ reasoning and tool-use capabilities to more closely simulate human cognitive processes in fact-checking.
3 Framework
Assessing the factual accuracy of long-form text presents significant challenges (Min et al., 2023). To address this, prior approaches have broken down the text into individual checkworthy claims (Chern et al., 2023). These sentences, referred to as atomic claims, are fact-checked individually, with their factuality scores aggregated to evaluate the overall factual accuracy of the original text. Previous research indicates that verifying the factuality of atomic claims is the most challenging step in this process (Wang et al., 2024a). Our work therefore focuses on this critical task: determining the factual accuracy of individual atomic claims, classifying each as either True or False.
3.1 Fire
We present Fire, a simple yet effective agent-based framework for interactive claim verification through web searches. As illustrated in Figure 1, Fire takes an atomic claim as input and outputs a binary label indicating whether the claim is factual or non-factual. The framework consists of three key components: Final Answer or Next Search Query, Web Search, and Final Verification, each of which we will explain below.
Final Answer or Next Search Query
We introduce a unified method, Final Answer or Next Search Query $f(\cdot)$, which integrates claim verification with search query generation. Given an atomic claim $c$, this component decides whether to produce a final answer $a$ or generate an additional search query $q$. This decision is guided by both an external evidence set $E$, derived from search results, and the internal knowledge $k$ of the language model, acquired during pre-training. At the outset, no evidence has been retrieved, meaning that the evidence set $E$ is initially empty. Consequently, the decision relies solely on the internal knowledge $k$. As shown in Equation 1, we incorporate confidence estimation into the reasoning process to determine the next action. If the model’s confidence is sufficiently high, it outputs a final answer $a$; otherwise, it generates an additional query $q$.
$f(c,E,k)=\begin{cases}a,&\text{ if confident}\\ q,&\text{ if not confident}\end{cases}$ (1)
This method offers greater flexibility by eliminating the need to retrieve a fixed number of evidence items before verification, thereby largely reducing search costs. A detailed description of the prompt used for this component is provided in Appendix A.
Web search
When the language model determines that a web search is necessary and issues a search query $q$, we retrieve results using Google Search via the SerpAPI [^1] https://serpapi.com, following prior work (Wei et al., 2024). This API returns the retrieved snippets as a single string, which we use as the new evidence $e$. We then append $e$ to the existing evidence set $E$ to form the updated evidence set $E^{\prime}$ for the next iteration, as shown in Equation 2.
$E^{\prime}=E\cup e,e=\text{Search}(q)$ (2)
Final Verification
Due to the inherent difficulty of confidently verifying certain claims, even with supplementary evidence, we impose an upper limit on the number of retrieval steps. As shown in Equation 3, once this limit is reached, the model performs a final verification $f^{\prime}(\cdot)$ based on all previously retrieved evidence. The detailed prompt for this process is provided in Appendix B.
$\begin{cases}a=f^{\prime}(c,E,k),&n\geq N\\ e=\text{Search}(q),&n<N\end{cases}$ (3)
3.2 Prevention of Repetitive Search Queries
In our preliminary studies, we identified a recurring issue with sequential search query generation using language models: the tendency of these models to generate repetitive queries. This occurs even when the models are explicitly instructed to generate queries targeting new, claim-relevant information. As a result, identical queries are repeatedly submitted to web search tools, leading to inefficient use of search resources. To address this issue, we investigate following methods for enhancing search query generation and reducing repetition.
Early Termination
The iterative process is terminated when consecutive queries or retrieved results exhibit a high degree of similarity, indicating diminishing returns.
Diversity Prompt
We introduce additional prompts to encourage the model to generate more diverse queries when consecutive similar queries or search results are detected.
3.3 Prevention of Verification Overconfidence
LLMs can exhibit strong calibration abilities across diverse tasks (Kadavath et al., 2022; Geng et al., 2024). Consequently, they are aware of their confidence levels during the claim verification process. However, our preliminary analysis reveals that LLMs often demonstrate excessive strictness and unwarranted confidence in certain cases, leading to errors. Considering this, we explore several techniques to prevent overconfidence in verification:
At Least One/Two
At Least One requires models to retrieve at least one evidence during the verification, which increase the probability of eliminating overconfidence. Similarly, we also adopted a more aggressive approach At Least Two to retrieve a second evidence to reduce the uncertainty.
Inclusive Prompt
In this setting, we prompt models to be “less strict, open-minded and avoid being over confident” to encourage models to reflect on their confidence level of answers.
4 Experiments Setup
4.1 Datasets
In our study, we utilized four datasets from prior research that align with our experimental setup: FacTool (Chern et al., 2023), FELM (Chen et al., 2023), Factcheck-Bench (Wang et al., 2024a), and BingCheck (Li et al., 2024b). FacTool and FELM provide factuality claims across multiple domains. From these, we selected instances requiring world knowledge for verification, which we refer to as FacTool-QA and FELM-WK, both annotated with binary labels (True or False). Our selection was motivated by the need to focus on claims that challenge models to use external knowledge, a critical aspect of factual verification.
For Factcheck-Bench and BingCheck, we consolidated the original four-label classification (supported, partially supported, not supported, refuted) into a binary format by merging supported and partially supported into True, treating refuted as False, and excluding not supported. This binarization aligns these datasets with the others and simplifies evaluation, focusing on clear-cut factuality decisions. We sampled a subset of BingCheck due to its class imbalance (3,581 True claims versus 42 False claims), selecting 100 True claims for our test set. This sampling was essential to create a more balanced and manageable test set, ensuring that evaluation metrics reflect performance on both classes without being dominated by the majority class. In FELM-WK, we retained un-split claims to maintain contextual integrity, which is crucial for accurate verification. Full dataset statistics are provided in Table 1.
Dataset #True #False Total Factcheck-Bench 472 159 631 FacTool-QA 177 56 233 FELM-WK 99 85 184 BingCheck 100 42 142
Table 1: Statistics of the datasets after processing.
In our experiments, we first use the Factcheck-Bench dataset as a development set to optimize the settings for our framework. We then evaluate its performance on the remaining three datasets, comparing it with other competitive fact-checking systems.
4.2 Language Models
We investigate several state-of-the-art (SOTA) language models, including proprietary models from two prominent families: GPT models (OpenAI, 2024a, b) and Claude models (Anthropic, 2024), as detailed in Table 2. In addition, we assess two open-source models: LLaMA 3.1-Inst 8B (Dubey et al., 2024) and Mistral-Inst 7B (Jiang et al., 2023).
Family Name
GPT GPT-4o, GPT-4o-mini, o1-preview, o1-mini
Claude Claude-3 Haiku, Claude-3 Opus, Claude-3.5 Sonnet
LLaMA LLaMA 3.1-Inst 8B
Mistral Mistral-Inst 7B
Table 2: Model families and specific model names used in this study.
4.3 Compared Fact-checking Frameworks
We select several SOTA fact-checking frameworks for comparison. Additionally, we introduce two baseline models: Random and Always True/False. To further assess the impact of LLM reasoning and evidence retrieval in fact-checking, we include two ablation settings: Fire (No Reason) and Fire (No Search).
FacTool
is adaptable across domains and tasks, using a tool-augmented framework that integrates external tools like Google Search and Python interpreters to assess the factuality of content from large language models. However, this can introduce complexity and depend on the accuracy of these external tools.
Factcheck-GPT
excels in fine-grained factuality evaluation through a detailed benchmark with annotations at the claim, sentence, and document levels. While resource-intensive, it provides valuable insights into specific stages of factual inaccuracies.
Safe
uses a search-augmented approach to verify long-form content by breaking it down into individual facts and checking them via Google Search. This method is cost-effective compared to human annotation but depends on the reliability of search engine results, which can vary and introduce biases.
Random
assigns the predicted label for each claim in the test set randomly, choosing between True and False with equal probability.
Always True/False
is an approach that always predicts a single label – either True or False – for all claims in the test set.
Fire (No Reason)
utilizes the same framework as Fire; however, it is explicitly instructed not to articulate its reasoning process in the output. This modification aims to assess the impact of explicitly stating the step-by-step reasoning process on the results.
Fire (No Search)
employs the same framework as Fire; however, it is not permitted to invoke the search tool. This configuration is designed to evaluate the model’s ability to perform fact-checking without retrieving any supporting evidence.
4.4 Evaluation Metrics
In this work, we investigate the trade-off between computational cost and fact-checking performance.
Performance
We evaluate precision, recall, and F1 scores for both positive and negative classes.
Computational Cost
We report the financial costs of LLM API calls for proprietary models and GPU rental expenses for open-source models, alongside an analysis of API costs from search engine queries and a breakdown of the total time spent on the fact-checking process. The experiments using open-source models were conducted on an NVIDIA RTX 6000 GPU at an estimated cost of \$0.79 per hour, while search queries via SerpAPI incurred approximately \$0.00105 per search.
5 Results
In this section, we first present preliminary studies on Factcheck-Bench (§ 5.1), focusing on three key aspects: language models, prevention of repetitive search queries, and prevention of verification overconfidence. These studies aim to identify the most appropriate configurations for our framework. Subsequently, we compare Fire to other strong fact-checking frameworks across three additional datasets (§ 5.2) to evaluate the generalization capabilities of our approach.
5.1 Preliminary studies
Language Models
We present a performance comparison of various language models in Table 3. Overall, proprietary language models generally outperform open-source models, likely due to their larger size and more sophisticated training in reasoning and tool utilization. Among the proprietary models, the latest and most advanced offerings from different organizations—specifically o1-preview from OpenAI and Claude-3.5 Sonnet from Anthropic—exhibit the best performance. Although the more economical model, GPT-4o-mini, performs slightly worse than the top-performing o1-preview, it offers a cost savings of 766 times. This suggests that for fact-checking tasks, the most advanced models may not be necessary; GPT-4o-mini can serve as a sufficiently capable alternative at a significantly lower cost. We will continue our preliminary studies using GPT-4o-mini.
LLM LLM+Search Cost (\$) Label = True Label = False
Prec Recall F1 Prec Recall F1
GPT-4o-mini 0.19+0.44 0.91 0.84 0.87 0.61 0.74 0.67
GPT-4o 10.45+1.47 0.92 0.79 0.85 0.56 0.79 0.66
o1-preview 145.66+0.80 0.91 0.86 0.88 0.64 0.75 0.69
o1-mini 20.06+1.13 0.89 0.81 0.85 0.56 0.71 0.62
Claude-3 Haiku 0.56+0.85 0.9 0.81 0.85 0.56 0.73 0.64
Claude-3 Opus 48.64+1.43 0.92 0.81 0.86 0.58 0.79 0.67
Claude-3.5 Sonnet 13.21+1.63 0.94 0.79 0.86 0.58 0.85 0.69
LLaMA 3.1-Inst 8B 3.95+2.27 0.89 0.74 0.8 0.48 0.72 0.57
Mistral-Inst 7B 1.84+1.22 0.85 0.67 0.75 0.4 0.66 0.5
Table 3: Fact-checking performance and cost comparisons between different language models within Fire on Factcheck-Bench.
Prevention of Repetitive Search Queries
We conducted an experimental analysis to evaluate the impact of Early Termination and Diversity Prompt on mitigating the generation of repetitive search queries. To assess query similarity, we employed Sentence-BERT (all-MiniLM-L6-v2; Reimers and Gurevych (2019)) with a similarity threshold of 0.9, as established by Shashavali et al. (2019). Table 4 presents experimental results, where window size refers to the predefined number of consecutive similar queries or retrieval results. Once this threshold is reached, early termination is triggered to prevent further query generation and retrieval. If the model generates queries or retrieves results exhibiting high similarity within this window, the system also activates an early stopping mechanism. The results indicate that optimizing the similarity window size effectively reduces search costs without compromising the model’s performance. However, our findings suggest that the diversity prompt does not enhance performance. In our optimal configuration, we selected a window size of 2 without utilizing the diversity prompt.
Window Size Diversity Prompt LLM+Search Cost (\$) Label = True Label = False
Prec Recall F1 Prec Recall F1
2 0.17+0.29 0.92 0.83 0.87 0.61 0.77 0.68
0.16+0.29 0.91 0.81 0.86 0.57 0.76 0.65
3 0.17+0.36 0.91 0.82 0.87 0.60 0.77 0.67
0.18+0.36 0.91 0.82 0.86 0.59 0.76 0.66
4 0.18+0.39 0.91 0.81 0.86 0.57 0.76 0.65
0.18+0.39 0.91 0.82 0.86 0.59 0.76 0.66
Default - 0.19+0.44 0.91 0.84 0.87 0.61 0.74 0.67
Table 4: Fire performance across various window sizes, with and without the use of prompts for generating diverse queries on Factcheck-Bench.
Prevention of Verification Overconfidence
We present the performance and cost of various overconfidence prevention approaches for verification on Factcheck-Bench in Table 5. Interestingly, the At Least One/Two settings, which aggressively retrieve additional evidence, result in higher search costs without improving fact-checking performance compared to the Default setting, where no explicit constraints are placed on web search. This supports our hypothesis that most atomic claims are relatively straightforward and do not require extensive external web searches for verification. In fact, introducing additional searches may introduce noise, negatively impacting performance. The Inclusive setting encourages models to be more flexible and open to alternative interpretations of evidence, which reduces the need for queries but also leads to lower overall performance. Based on these observations, we maintain the Default setting, leveraging the language model’s reasoning capabilities without imposing additional search constraints.
Approach LLM+Search Cost (\$) Label = True Label = False
Prec Recall F1 Prec Recall F1
At Least One 0.21+0.83 0.92 0.81 0.86 0.58 0.78 0.67
At Least Two 0.22+0.87 0.91 0.79 0.84 0.55 0.77 0.64
Inclusive 0.20+0.42 0.91 0.81 0.86 0.58 0.77 0.66
Default 0.19+0.44 0.91 0.84 0.87 0.61 0.74 0.67
Table 5: Fire performance using different verification overconfidence prevention approaches on Factcheck-Bench.
Framework LLM FacTool-QA FELM-WK BingCheck
Label = True Label = False Label = True Label = False Label = True Label = False
Prec Recall F1 Prec Recall F1 Prec Recall F1 Prec Recall F1 Prec Recall F1 Prec Recall F1
Random - 0.81 0.47 0.59 0.28 0.64 0.39 0.75 0.49 0.59 0.30 0.57 0.39 0.77 0.67 0.72 0.40 0.52 0.45
Always True - 0.76 1.0 0.86 0 0 0 0.72 1.0 0.84 0 0 0 0.70 1.0 0.83 0 0 0
Always False - 0 0 0 0.24 1.0 0.39 0 0 0 0.28 1.0 0.44 0 0 0 0.30 1.0 0.46
FacTool GPT-4o 0.88 0.81 0.84 0.52 0.66 0.58 0.69 0.53 0.60 0.57 0.73 0.64 0.86 0.57 0.68 0.43 0.79 0.56
GPT-4o-mini 0.92 0.68 0.78 0.45 0.82 0.58 0.67 0.37 0.48 0.51 0.78 0.62 0.92 0.55 0.69 0.45 0.88 0.60
Factcheck-GPT GPT-4o 0.90 0.79 0.84 0.52 0.71 0.60 0.67 0.68 0.67 0.61 0.61 0.61 0.85 0.70 0.77 0.50 0.71 0.59
GPT-4o-mini 0.85 0.80 0.82 0.47 0.56 0.51 0.61 0.50 0.55 0.51 0.62 0.56 0.88 0.78 0.83 0.60 0.76 0.67
Safe GPT-4o 0.92 0.88 0.90 0.66 0.77 0.71 0.70 0.80 0.75 0.72 0.60 0.65 0.84 0.90 0.87 0.71 0.60 0.65
GPT-4o-mini 0.92 0.82 0.87 0.58 0.79 0.67 0.61 0.76 0.68 0.61 0.44 0.51 0.86 0.81 0.84 0.60 0.69 0.64
Fire GPT-4o 0.92 0.88 0.90 0.65 0.71 0.68 0.70 0.86 0.77 0.77 0.54 0.63 0.86 0.88 0.87 0.70 0.67 0.68
GPT-4o-mini 0.87 0.88 0.87 0.60 0.59 0.59 0.63 0.82 0.71 0.67 0.44 0.53 0.87 0.91 0.88 0.74 0.67 0.70
Fire (No Reason) GPT-4o 0.88 0.86 0.87 0.60 0.64 0.62 0.70 0.85 0.77 0.77 0.58 0.66 0.85 0.89 0.87 0.70 0.62 0.66
GPT-4o-mini 0.87 0.84 0.86 0.55 0.61 0.58 0.65 0.84 0.73 0.71 0.47 0.57 0.84 0.87 0.85 0.66 0.6 0.62
Fire (No Search) GPT-4o 0.86 0.87 0.88 0.61 0.54 0.57 0.69 0.86 0.77 0.77 0.55 0.65 0.86 0.91 0.88 0.79 0.64 0.71
GPT-4o-mini 0.84 0.84 0.84 0.49 0.48 0.49 0.61 0.86 0.72 0.7 0.36 0.48 0.83 0.9 0.87 0.71 0.57 0.63
Table 6: Performance comparisons between different frameworks across multiple datasets.
Framework LLM LLM Search Time
FacTool GPT-4o 24.76 3.67 2.92
GPT-4o-mini 1.49 3.67 2.34
Factcheck-GPT GPT-4o 21.41 - 4.25
GPT-4o-mini 1.28 - 4.09
Safe GPT-4o 6.34 2.93 4.62
GPT-4o-mini 0.43 2.93 4.25
Fire GPT-4o 3.35 0.60 1.31
GPT-4o-mini 0.14 0.20 1.25
Fire (No Reason) GPT-4o 1.65 0.68 0.57
GPT-4o-mini 0.07 0.59 0.54
Fire (No Search) GPT-4o 1.70 - 1.03
GPT-4o-mini 0.11 - 1.34
Table 7: LLM/Search cost (USD) and time (hrs) for evaluating the total 559 atomic claims in FacTool-QA, FELM-WK, and BingCheck. We use SerperAPI for FacTool, Safe and Fire for search, while Factcheck-GPT has its own implemented scrapping technique.
5.2 Comparisons to Other Frameworks
We present a performance comparison of our framework against other frameworks in Table 6 and a cost analysis in Table 7. As shown, all frameworks exhibit similar performance, with a small gap of approximately 0.2. Our framework, using GPT-4o, performs slightly better, achieving superior results on 7 out of 18 metrics, followed closely by Safe with GPT-4o at 6 metrics. This suggests that all frameworks can effectively perform fact-checking for most claims, although they may encounter difficulties with challenging examples, which we analyze further in § 6. Regarding the necessity of evidence retrieval in fact-checking, we observe a relatively larger performance drop in FacTool when evidence search is omitted, compared to a smaller drop in FELM-WK and BingCheck. This suggests that FacTool-QA comprises more rare knowledge than GPT-4o-mini, whereas FELM-WK and BingCheck may rely predominantly on common knowledge, for which evidence retrieval is less impactful. Overall, both GPT-4o and GPT-4o-mini perform reasonably well on popular public datasets, highlighting the need for datasets that incorporate more complex claims. In terms of model size, GPT-4o generally outperforms GPT-4o-mini across most frameworks, indicating that larger models are more effective in detecting misinformation. However, the performance improvement is limited, and the associated costs result in an average increase of 16.7 times in LLM expenses and a three-fold increase in search costs when using Fire. Therefore, we argue that cheaper models, such as GPT-4o-mini, are a viable option for performing fact-checking tasks. Furthermore, when considering all frameworks with GPT-4o-mini, Fire achieves additional cost savings, reducing LLM expenses by 7.6 times and search costs by 16.5 times compared to other frameworks. Thus, we contend that Fire, when paired with GPT-4o-mini, offers a compelling solution for the large-scale deployment of fact-checking systems.
Figure 2 illustrates the impact of reasoning on the number of web searches conducted by GPT-4o and GPT-4o-mini tested on BingCheck. Notably, GPT-4o-mini demonstrates a high level of confidence in making verifications when it is allowed to articulate its reasoning process, resulting in the majority of judgments being made without any searches. Conversely, when not permitted to express its reasoning, there is a significant decrease in the number of instances with zero searches; most cases now involve at least one search, indicating a marked reduction in GPT-4o-mini’s confidence in its judgments. This observation aligns with previous findings that the presence of CoT reasoning correlates with increased confidence in the model’s answers (Wang and Zhou, 2024). While GPT-4o also shows a decline in confidence when it is not allowed to search, the decrease is less pronounced than that observed in GPT-4o-mini.
By combining the performance and cost results presented in Table 6 and Table 7, we find that, in the absence of a reasoning process, the costs associated with LLMs can be reduced through fewer completion tokens. However, this reduction leads to increased search costs, resulting in overall performance that is inferior to scenarios in which the models are permitted to engage in step-by-step reasoning. Furthermore, the step-by-step reasoning approach facilitates more effective error analysis.
Figure 2: The effect of reasoning on the number of searches using GPT-4o and GPT-4o-mini within Fire on BingCheck. The shaded area indicates the number of misclassified cases. The x-axis shows the number of web searches, while the y-axis denotes the number of instances.
Major Issue Error Type Description FELM FacTool BingCheck #Total
I. Dataset Issue 1. Not a claim, e.g. a claim only has a name Elvis Presley. 12 1 0 13
2. Unclear, ambiguous or subjective claim e.g. there is no record of how many sons he had. 11 7 2 20
3. False gold labels, i.e., the original annotated label might be wrong. For example, claim choosing organic and local foods that are in season can reduce emissions from transporting food from far away is labeled as false 10 0 1 11
II. Knowledge Issue 4. Complicated science domain expert knowledge is needed to judge, like astronomy. 3 0 1 4
5. Inaccurate parametric knowledge. LLM-based verifiers make wrong verification due to the incorrect parametric knowledge stored in LLMs. 3 6 7 16
6. Insufficient or inaccurate externally collected knowledge (evidence), involving three scenarios: (i) no external evidence, model makes wrong reasoning by itself; (ii) collected evidence is insufficient to cover all aspects mentioned by a long claim; (iii) collected evidence is inaccurate (e.g., evidence contain statement More than 430 species of mammal are found in the Amazon when the correct number is 427). 9 15 6 30
III. LLM Reasoning 7. Incorrect reasoning, e.g., the claim mentioned A while the model dismissed in the reasoning process, or the claim did not mention A while the model hallucinated A 6 8 4 18
8. Strict reasoning includes two situations: (i) strictly depending on the collect evidence to make decision leads to wrong verification, while if it combines commonsense and collected evidence to analyze, it can verify correctly; (2) strict reasoning based on parametric knowledge. E.g. regarding Word-Cross Puzzle and FUN Word-Cross Puzzle are different. 5 0 3 8
IV. Debatable Opinion 9. Debatable Opinions on controversial topics, e.g. actual origins are debated for claim Fortune cookies made their way to San Francisco in the late 1800s and early 1900s through Japanese immigrants. 7 8 0 15
Total 66 45 24 135
Table 8: Datasets Error distribution, grouped into nine fine-grained types under four major issues.
6 Error Analysis
To identify weaknesses in our fact-checking system, we manually examine failed cases of three datasets: FELM-WK, FacTool-QA, and BingCheck, analyzing whether the majority of failures is attributed to inadequate retrieved evidence or to flaws in the LLM verification process, despite the availability of reliable evidence.
We summarized errors into four major issues and nine error types. Among the total number of 135 failed claims, there are 44 cases falling into challenges of (I) inaccurate identification of checkworthy claims and false gold labels in the original datasets, 50 claims are due to (II) inaccurate or insufficient knowledge applied to verification, either internally extracted from LLM parameters or externally collected from web pages. The rest 26 and 15 cases result from LLM reasoning ability and debatable opinions over some topics, respectively, as shown in  Table 8.
The major issue lies in collecting sufficient evidence, especially for long claims containing many aspects to verify. This can be approached by decomposing “atomic claims” from the original dataset into the real granularity of “atomic”, each containing only 1-3 pieces of information. The second problem focus on the quality of benchmarking datasets, particularly FELM-WK that includes many ungrounded claims and labels (Li et al., 2024a), which may lead to ineffective comparisons between fact-checking systems. Interestingly, beyond incorrect reasoning, overly-strict reasoning by exact matching between the claim and collected evidence can also lead to verification errors. For example, LLMs label a claim as false when the claim states FUN Word-Cross Puzzle while evidence mentions Word-Cross Puzzle. Additionally, some claims can be viewed as true from one perspective but false from another, as seen in debates over the origins of fortune cookies, where the truth of related claims is debatable.
Considering above, to further advance the field of fact-checking, we highlight the need for improved benchmarking datasets, a stronger focus on verifying fine-grained claims, and strategies to guide LLMs in performing verification under more flexible reasoning conditions, such as semantic alignment, rather than relying exclusively on exact matches.
7 Conclusions and Future Work
Conventional fact-checking systems typically separate the steps of evidence retrieval and claim verification, leading to suboptimal utilization of the verification models’ internal knowledge. To address this, we propose Fire, a novel framework that integrates evidence retrieval and claim verification in an iterative process. Fire enables LLMs to leverage their internal knowledge for judgment and only rely on external evidence retrieval when uncertain. Our experiments on multiple datasets demonstrate that Fire not only slightly improves accuracy but also reduces LLM computation costs by an average of 7.6 times and search costs by 16.5 times, making it highly efficient for production use. Additionally, we performed a detailed error analysis, which revealed issues with the benchmarking datasets quality. These findings highlight the need for further research into edge cases, rather than relying solely on automatic metrics for evaluation.
We identify several promising directions for future work, which include: (1) Integrating memory banks to store verification results, allowing the system to reuse previous results instead of repeatedly executing the entire process; (2) Expanding the system to support additional modalities, such as code and images; and (3) Revisiting existing public fact-checking datasets, incorporating personal opinions when addressing ambiguous cases, and adding claims that require rarer and more complex knowledge, where evidence retrieval is essential.
Limitations
We acknowledge several limitations in this work that we plan to address in future research. First, to maintain the efficiency of our framework, we implement the “Final Answer or Next Search Query” mechanism in a compact manner, allowing it to retrieve evidence, assess confidence in knowledge, and verify the final answer within a single step. Ideally, this process could be separated to include a standalone confidence estimation step, which would enhance both flexibility and interpretability. We leave this exploration to future work. Second, to ensure a fair comparison across multiple fact-checking datasets, our system adopts a binary labeling scheme (“True” or “False”) and standardizes labels across datasets. However, this approach may not fully capture the complexity of factual labels in real-world settings. We intend to incorporate fine-grained labeling schemes in future research. Finally, in this study, we rely on SerpAPI with its default settings. While we did not investigate in detail how evidence is retrieved, we believe future work could explore this aspect further to optimize the selection of the most relevant evidence for a given claim.
Ethical Statement and Broad Impact
Data License
A primary ethical consideration is the data license. We reused pre-existing dataset, FactBench, FacTool, FELM-WK, BingCheck, which have been publicly released and approved for research purposes. We adhere to the intended usage of all these dataset licenses.
Ethical Statement
We acknowledge that our system relies on LLMs, which can sometimes produce biased or incorrect judgments due to the data used in their pre-training or biases present in external sources. Additionally, there is the risk of over-reliance on the system for making critical factual judgments without human oversight. To mitigate these risks, we strongly encourage human reviewers to be involved in decision-making, especially in high-stakes domains such as legal, political, or medical contexts.
Broad Impact
Fire has the potential to advance the field of automated fact-checking by enhancing its efficiency and accessibility. Its capability to iteratively retrieve evidence while minimizing computational costs will empower a broader range of users—including journalists, researchers, and the general public—to verify factual information with greater ease. Furthermore, Fire can be applied to large-scale implementations, such as integration into search engines and social media platforms, thereby contributing to efforts to combat the spread of misinformation.
Acknowledgments
We thank our reviewers for their valuable reviews and feedback, which significantly contributed to the improvement of our paper.
Appendix A Prompts for Verification
Default prompt
We use the following prompt to guide the language model in verifying the atomic claim, determining whether to provide a final judgment or issue an additional Google search query based on the current status. The prompt will output reason or explanation for the verification process.
_FINAL_ANSWER_OR_NEXT_SEARCH_FORMAT = f"""Instructions:
1. You are provided with a STATEMENT and relevant KNOWLEDGE points.
2. Based on the KNOWLEDGE, assess the factual accuracy of the STATEMENT.
3. Before presenting your conclusion, think through the process step-by-step. Include a summary of the key points from the KNOWLEDGE as part of your reasoning.
4. If the KNOWLEDGE allows you to confidently make a decision, output the final answer as a JSON object in the following format:
"final_answer": "{_Factual_LABEL}" or "{_Non_Factual_LABEL}"
5. If the KNOWLEDGE is insufficient to make a judgment, issue ONE Google Search query that could provide additional evidence. Output the search query in JSON format, as follows:
"search_query": "Your Google search query here"
6. The query should aim to obtain new information not already present in the KNOWLEDGE, specifically helpful for verifying the STATEMENTs accuracy.
KNOWLEDGE:
{_KNOWLEDGE_PLACEHOLDER}
STATEMENT:
{_STATEMENT_PLACEHOLDER}
No Reason prompt
To improve efficiency, we opted for this setting to output only the label.
_FINAL_ANSWER_OR_NEXT_SEARCH_FORMAT = f"""Instructions:
1. You are provided with a STATEMENT and relevant KNOWLEDGE points.
2. Based on the KNOWLEDGE, assess the factual accuracy of the STATEMENT.
3. Before presenting your conclusion, think through the process step-by-step. Include a summary of the key points from the KNOWLEDGE as part of your reasoning.
4. If the KNOWLEDGE allows you to confidently make a decision, output the final answer as a JSON object in the following format:
"final_answer": "{_Factual_LABEL}" or "{_Non_Factual_LABEL}"
5. If the KNOWLEDGE is insufficient to make a judgment, issue ONE Google Search query that could provide additional evidence. Output the search query in JSON format, as follows:
"search_query": "Your Google search query here"
6. The query should aim to obtain new information not already present in the KNOWLEDGE, specifically helpful for verifying the STATEMENTs accuracy.
7. Do not provide any additional information or reasoning in the output. Only output the JSON object.
KNOWLEDGE:
{_KNOWLEDGE_PLACEHOLDER}
STATEMENT:
{_STATEMENT_PLACEHOLDER}
At Least One prompt
At Least One prompt requires models to retrieve at least one evidence during the verification.
_FINAL_ANSWER_OR_NEXT_SEARCH_FORMAT = f"""Instructions:
1. You are provided with a STATEMENT and relevant KNOWLEDGE points.
2. Based on the KNOWLEDGE, assess the factual accuracy of the STATEMENT.
3. Before presenting your conclusion, think through the process step-by-step. Include a summary of the key points from the KNOWLEDGE as part of your reasoning.
4. If the KNOWLEDGE allows you to confidently make a decision, output the final answer as a JSON object in the following format:
"final_answer": "{_Factual_LABEL}" or "{_Non_Factual_LABEL}"
5. If the KNOWLEDGE is insufficient to make a judgment, issue ONE Google Search query that could provide additional evidence. Output the search query in JSON format, as follows:
"search_query": "Your Google search query here"
6. The query should aim to obtain new information not already present in the KNOWLEDGE, specifically helpful for verifying the STATEMENTs accuracy.
7. If the KNOWLEDGE is empty, please issue ONE Google Search query immediately.
KNOWLEDGE:
{_KNOWLEDGE_PLACEHOLDER}
STATEMENT:
{_STATEMENT_PLACEHOLDER}
At Least Two prompt
At Least Two prompt is a more aggressive approach to retrieve minimum two evidence before verification.
_FINAL_ANSWER_OR_NEXT_SEARCH_FORMAT = f"""Instructions:
1. You are provided with a STATEMENT and relevant KNOWLEDGE points.
2. Based on the KNOWLEDGE, assess the factual accuracy of the STATEMENT.
3. Before presenting your conclusion, think through the process step-by-step. Include a summary of the key points from the KNOWLEDGE as part of your reasoning.
4. If the KNOWLEDGE allows you to confidently make a decision, output the final answer as a JSON object in the following format:
"final_answer": "{_Factual_LABEL}" or "{_Non_Factual_LABEL}"
5. If the KNOWLEDGE is insufficient to make a judgment, issue ONE Google Search query that could provide additional evidence. Output the search query in JSON format, as follows:
"search_query": "Your Google search query here"
6. The query should aim to obtain new information not already present in the KNOWLEDGE, specifically helpful for verifying the STATEMENTs accuracy.
7. If the KNOWLEDGE is empty or there is only ONE evidence in the KNOWLEDGE, please issue ONE Google Search query immediately.
KNOWLEDGE:
{_KNOWLEDGE_PLACEHOLDER}
STATEMENT:
{_STATEMENT_PLACEHOLDER}
Inclusive
In this setting, we prompt models to be “less strict, open-minded and avoid being over confident” to encourage models to reflect on their confidence level of answers.
_FINAL_ANSWER_OR_NEXT_SEARCH_FORMAT = f"""Instructions:
1. You are provided with a STATEMENT and relevant KNOWLEDGE points.
2. Based on the KNOWLEDGE, assess the factual accuracy of the STATEMENT.
3. Before presenting your conclusion, think through the process step-by-step. Include a summary of the key points from the KNOWLEDGE as part of your reasoning.
4. If the KNOWLEDGE allows you to confidently make a decision, output the final answer as a JSON object in the following format:
"final_answer": "{_Factual_LABEL}" or "{_Non_Factual_LABEL}"
5. If the KNOWLEDGE is insufficient to make a judgment, issue ONE Google Search query that could provide additional evidence. Output the search query in JSON format, as follows:
"search_query": "Your Google search query here"
6. The query should aim to obtain new information not already present in the KNOWLEDGE, specifically helpful for verifying the STATEMENTs accuracy.
7. Please be more open-minded and less strict in your evaluation. Avoid being overly confident, and consider the possibility of alternative interpretations or uncertainties in the evidence.
KNOWLEDGE:
{_KNOWLEDGE_PLACEHOLDER}
STATEMENT:
{_STATEMENT_PLACEHOLDER}
Appendix B Prompt for Final Verification
Upon reaching the maximum number of steps, we issue the following prompt to compel the language model to make a final judgment based on the accumulated information.
_MUST_HAVE_FINAL_ANSWER_FORMAT = f"""Instructions:
1. You are provided with a STATEMENT and relevant KNOWLEDGE points.
2. Based on the KNOWLEDGE, assess the factual accuracy of the STATEMENT.
3. Before presenting your final answer, think step-by-step and show your reasoning. Include a summary of the key points from the KNOWLEDGE as part of your reasoning.
4. Your final answer should be either "{_Factual_LABEL}" or "{_Non_Factual_LABEL}".
5. Format your final answer as a JSON object in the following structure:
"final_answer": "{_Factual_LABEL}" or "{_Non_Factual_LABEL}"
6. Do not include any other information or reasoning in the output. Only provide the JSON object.
KNOWLEDGE:
{_KNOWLEDGE_PLACEHOLDER}
STATEMENT:
{_STATEMENT_PLACEHOLDER}
Appendix C Effect of Reasoning
We additionally include figures to illustrate the effect of reasoning on two other datasets: FacTool-QA (Figure 3) and FELM-WK (Figure 4), supplementing Figure 2. These figures demonstrate that both GPT-4o and GPT-4o-mini are influenced by explicitly stating their reasoning process, with GPT-4o-mini showing a consistent impact across all datasets, not just BingCheck. Furthermore, when comparing these datasets, we observe that the models appear most confident on FELM-WK compared to the other two datasets. As a result, even in the absence of explicit reasoning, they do not perform any searches to verify the claims.
Figure 3: The effect of reasoning on the number of searches using GPT-4o and GPT-4o-mini within Fire on FacTool-QA.
Figure 4: The effect of reasoning on the number of searches using GPT-4o and GPT-4o-mini within Fire on FELM-WK.